Consider the table. Labor Quantity (hours) 0 0 Variable Cost Total Cost Marginal Cost ($) Average Cost ($) AVC* ($) 0 40 18 5E: F: G: H:
Answer #1

A = 20000 because variable cost = total cost - fixed cost

So variable cost = 30000 - 10000 = 20000

B = 20000 because total cost = fixed cost + variable cost = 10000 + 10000 = 20000

C = 277.7 because MCn = TCn - TCn-1 / quantity

MC = 15000 - 10000/18 = 5000/18 = 277.7

D = 384.6 because MC = 30000-25000/78-65 = 5000/13 = 384.6

E = 833.3 because MC = 40000-35000/94-88 = 833.3

F = 833.3 because Average cost = Total cost / quantity = 15000 / 18 = 833.3

G = 384.6 because AC = 30000/78 = 384.6

H = 319.14 because AVC = VC/Q = 30000/94 = 319.14